The value Io of at transistor is 100 if the value of the emitter current is 10mA, then determine the value of the conductor and base current


1
Expert's answer
2021-05-12T16:50:58-0400

"I_e=10mA
,


\beta=100"


Base current "I_b=\dfrac{I_e}{\beta+1}=\dfrac{10}{100+1}=\dfrac{10}{101}=0.099mA"


Conductor current "I_c=\beta I_b=100\times 0.099=9.9mA"